Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.30 ounces $ 26.00 per ounce $ 59.80
Direct labor 0.50 hours $ 14.00 per hour 7.00
Variable manufacturing overhead 0.50 hours $ 3.40 per hour 1.70
Total standard cost per unit $ 68.50

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 12,500 ounces at a cost of $305,625.
  2. There was no beginning inventory of materials; however, at the end of the month, 2,800 ounces of material remained in ending inventory.

  3. The company employs 21 lab technicians to work on the production of Fludex. During November, they each worked an average of 150 hours at an average pay rate of $12.00 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $4,200.

  5. During November, the company produced 4,200 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 21 technicians employed in the production of Fludex consisted of 4 senior technicians and 17 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1
Actual materials used 9700 =12500-2800
Actual labor hours 3150 =21*150
1
a
Materials price variance 19375 F =305625-(12500*26)
Materials quantity variance 1040 U =26*(9700-4200*2.3)
b
Yes, as materials variance is favorable
2
a
Labor rate variance 6300 F =3150*(12-14)
Labor efficiency variance 14700 U =14*(3150-4200*0.5)
b
No, as labor variance is unfavorable
3
Variable overhead rate variance 6510 F =4200-(3150*3.4)
Variable overhead efficiency variance 3570 U =3.4*(3150-4200*0.5)